Two impedances Z1 and Z2 are connected in parallel. The first branch takes a leading current of 16 A and has a resistance of 5 Ω, while the second branch takes a lagging current at 0.8pf. The applied voltage is 100+j200 V and the total power is 5 kW. Find branch impedances, total circuit impedance, branch currents and total circuit current.


1
Expert's answer
2021-03-01T02:14:52-0500

I1 = 16 0°

V = 100 + j200

P = 5 kW

P = P1 + P2 = I12R1 + I22R2

5000 W = 162 x 5 + I22R2

I22R2 = 3720 W

p.f. = cos "\phi" = 0.8

"\phi" = 37°

Z1 = V/I1 = (100 + j200) / 16 0°

Z = Z1 + Z2